# DualSpeechLM: Towards Unified Speech Understanding and Generation via Dual Speech Token Modeling with Large Language Models

### v1

@@ Abstract
Extending pre-trained Large Language Models (LLMs)’s speech understanding or generation abilities by introducing various effective speech tokens has attracted great attention in the speech community. However, building a unified speech understanding and generation model still faces the following challenges: (1) Due to the huge modality gap between speech tokens and text tokens, extending text LLMs to unified speech LLMs relies on large-scale paired data for fine-tuning, and (2) Generation and understanding tasks prefer information at different levels, e.g., generation benefits from detailed acoustic features, while understanding favors high-level semantics. This divergence leads to difficult performance optimization in one unified model. To solve these challenges, in this paper, we present two key insights in speech tokenization and speech language modeling. Specifically, we first propose an Understanding-driven Speech Tokenizer (USTokenizer), which extracts high-level semantic information essential for accomplishing understanding tasks using text LLMs. In this way, USToken enjoys better modality commonality with text, which reduces the difficulty of modality alignment in adapting text LLMs to speech LLMs. Secondly, we present DualSpeechLM, a dual-token modeling framework that concurrently models USToken as input and acoustic token as output within a unified, end-to-end framework, seamlessly integrating speech understanding and generation capabilities. Furthermore, we propose a novel semantic supervision loss and a Chain-of-Condition (CoC) strategy to stabilize model training and enhance speech generation performance. Experimental results demonstrate that our proposed approach effectively fosters a complementary relationship between understanding and generation tasks, highlighting the promising strategy of mutually enhancing both tasks in one unified model.111 Code and demo: https://github.com/lavendery/Unified-Understanding-and-Generalization.
@@ Introduction
Recent advancements in autoregressive large language models (LLMs) have demonstrated excellent performance in the natural language processing community (Achiam et al. 2023; Touvron et al. 2023; Team et al. 2023). Leveraging the powerful foundations of text LLMs, recent advancements have led to emergence of speech LLMs that possess speech understanding and generation capabilities. In addition to fine-tuning textual LLMs to separately perform speech understanding tasks (Gong et al. 2023; Tang et al. 2023; Chu et al. 2023; Wang et al. 2024a) and generation tasks (Wang et al. 2023; Anastassiou et al. 2024; Kim et al. 2024; Wang et al. 2024b; Yang et al. 2024; Jia et al. 2025), developing unified models that excel at both capabilities has been explored in recent years (Zhang et al. 2023a; Xie et al. 2024; Fu et al. 2025; Nguyen et al. 2025; Défossez et al. 2024; Xu et al. 2025). However, several limitations still remain.
@@
First, adapting pre-trained text LLMs to unified speech LLMs still relies heavily on large-scale paired speech-text data (Zhang et al. 2023a; Défossez et al. 2024; Xie et al. 2024; Xu et al. 2025). For example, SpeechGPT (Zhang et al. 2023a) and SpiritLM (Nguyen et al. 2025) require approximately 70K and 570K hours of paired data, respectively. This dependence stems from the substantial modality gap between speech and text, which hinders capability transfer. Second, existing speech LLMs struggle to meet the distinct informational needs of understanding and generation. As shown in Figure 1 (a) Left and (b) Left, the Baseline model—trained solely with acoustic tokens—exhibits a contradiction between tasks, i.e., improving one often leads to degradation of the other, highlighting its inability to balance both tasks effectively. In fact, generation tasks demand rich acoustic details (e.g., prosody, emotion, speaker traits) for high-fidelity synthesis (Zeghidour et al. 2021; Défossez et al. 2022; Kumar et al. 2023; Wang et al. 2023), which acoustic tokens capture well but lack high-level semantics (Shi et al. 2024; Dhawan et al. 2024; Anastassiou et al. 2024; Chang et al. 2024). Conversely, understanding tasks benefit from semantic features (Borsos et al. 2023; Rubenstein et al. 2023; Maiti et al. 2024), but semantic tokens inevitably compromise the acoustic details needed for natural speech generation.
@@
To solve these problems, we propose two key insights from the perspectives of speech tokenization and language modeling. First, we present an Understanding-driven Speech Tokenizer (USTokenizer) that can extract high-level semantic features, which are critical for understanding tasks. Unlike prior methods that rely on only self-supervised learning (SSL) representation quantization (Hsu et al. 2021; Chen et al. 2022) or automatic speech recognition (ASR)-based objectives (Du et al. 2024; Zeng et al. 2024) to capture semantics, our approach directly aligns speech tokenizer with the semantic understanding capabilities of text LLMs. This leads to USTokens that have inherently better alignment with text modality, significantly easing modality alignment when adapting text LLMs to speech LLMs. Secondly, building on USTokenizer, we introduce DualSpeechLM, a novel dual speech token modeling framework that effectively handles the distinct informational needs of understanding and generation within a unified end-to-end architecture. Unlike conventional approaches that use the same token as input and output of LLM, our model separates them by using USTokens as input and acoustic tokens as output. Specifically, the USTokenizer provides high-level semantic information to boost understanding tasks, while an AcousticGPT module restores fine-grained acoustic details, enabling diverse and realistic speech generation within a unified end-to-end framework. In addition, we introduce a semantic supervision loss and a Chain of Condition (CoC) strategy to stabilize training and further improve the performance of the unified framework. In summary, our contributions are as follows:
@@
The success of autoregressive language models (Achiam et al. 2023; Touvron et al. 2023; Team et al. 2023) has spurred progress in speech LLMs (Chu et al. 2023; Wang et al. 2024b; Défossez et al. 2024), where speech tokenizers are essential for converting continuous signals into discrete tokens (Yang et al. 2025). Speech tokenizers are typically categorized as acoustic or semantic (Borsos et al. 2023). Acoustic tokens, optimized for signal reconstruction (Zeghidour et al. 2021; Défossez et al. 2022; Wang et al. 2023), capture detailed acoustic features beneficial for generation, but perform poorly on understanding tasks like ASR (Shi et al. 2024; Dhawan et al. 2024; Anastassiou et al. 2024; Chang et al. 2024). Previous semantic tokenizers can be trained in two ways: (1) applying clustering (Borsos et al. 2023; Zhang et al. 2023a) or VQ (Huang, Meng, and Ko 2024) to representations of self-supervised learning (SSL) models  (Hsu et al. 2021; Chen et al. 2022). (2) applying a VQ layer to the intermediate layer of ASR models  (Du et al. 2024; Zeng et al. 2024). Although these semantic tokenizers have shown benefits for understanding tasks (Borsos et al. 2023; Rubenstein et al. 2023; Maiti et al. 2024), they do not explicitly consider the alignment to the modality with text-LLM. Furthermore, multi-codebook designs (Zhang et al. 2023b; Défossez et al. 2024; Yang et al. 2025) aim to capture semantics and acoustics in different codebooks jointly, but often introduce complexity when integrated with LLMs. In this work, we present a single VQ-codebook USTokenizer, which not only incorporates high-level semantic information but also achieves better alignment between speech tokens and text modality when applied to LLMs.
@@
Recent advances in speech LLMs have explored unified understanding and generation (Zhang et al. 2023a; Pan et al. 2024; Défossez et al. 2024; Nguyen et al. 2025). Some speech LLMs (Yang et al. 2024; Shi et al. 2025) adopt acoustic tokens to ensure high-fidelity speech synthesis. However, such tokens usually lead to degraded performance in understanding tasks, particularly in low-resource scenarios. By contrast, semantic tokens perform better in the understanding task. Yet, their lack of acoustic details often results in reduced generation quality. To compensate for generation, existing works (Polyak et al. 2021; Zhang et al. 2023a; Du et al. 2024; Nguyen et al. 2025) introduce additional components such as diffusion models (Ho, Jain, and Abbeel 2020) or flow matching (Lipman et al. 2022), to convert speech tokens into a Mel spectrogram, and then a HiFi-GAN (Kong, Kim, and Bae 2020) vocoder is used to synthesize waveform with the Mel spectrogram as input. These multi-stage pipelines increase complexity and risk error accumulation (Jia et al. 2025). To address these limitations, we propose DualSpeechLM, an end-to-end dual-token modeling framework that explicitly models USTokens as input for understanding and acoustic tokens as output for generation, effectively achieving distinct informational requirements.
@@
The Encoder converts Whisper features $\mathbf{E}$ into representations $\mathbf{H}$ using a 2× downsampling convolution followed by two residual convolutional blocks, then discretized by VQ. Symmetrically, the Decoder reconstructs $\mathbf{\hat{E}}$ from quantized vectors $\mathbf{\hat{H}}$. Both Encoder and Decoder adopt similar residual convolutional blocks with RepCodec (Huang, Meng, and Ko 2024). We then compute the reconstruction loss via mean squared error (MSE) between $\mathbf{E}$ and $\mathbf{\hat{E}}$, formulated as:
@@
Figure 3: DualSpeechLM’s dual-token modeling paradigm. The left illustrates the baseline pipeline treating LLM input/output as identical tokens. In contrast, DualSpeechLM incorporates an Acoustic GPT module into the text LLM module for joint training, separately processing USToken inputs and acoustic token outputs through distinct modeling paths, effectively capturing the different levels of information required for both generation and understanding tasks.
@@
SpeechGPT  (Zhang et al. 2023a) LLaMA-7B D(Hubert) Full FT 70K 42.73 78.54 1.07 - 3.58/40
SpiritLM  (Nguyen et al. 2025) LLaMA-7B D(Hubert) Full FT 570K 6.0 11.0 - - -
Mini-Omni2  (Xie et al. 2024) Qwen2-0.5B C(Whisper) Full FT 9K 4.8 9.8 - - -
VITA  (Fu et al. 2025) Mixtral-8x7B C(Mel) Full FT 20K 8.14 18.41 - - 7.62/70
Moshi  (Défossez et al. 2024) Helium-7B D(Mimicodec) Full FT 7M 5.7 - - - -
Qwen2.5-Omni  (Xu et al. 2025) Qwen2.5-7B C(Whisper) Full FT $\star$ 1.8 3.4 30.2 - 4.28/60
@@
The VQ module discretizes continuous feature representations $\mathbf{H}$ by mapping them to a codebook of learned vectors. In this work, we use a single VQ layer to quantize the feature $\mathbf{H}$ into discrete tokens $\mathbf{Z}$. Following previous works (Huang, Meng, and Ko 2024), we add a commitment loss $\mathcal{L}_{\text{commit}}$ to ensure training stability, with more details provided in Appendix B.
@@
As shown in Figure 3, unlike prior Speech LLMs that use the same token for both input and output (Yang et al. 2024; Wang et al. 2024b; Shi et al. 2025), DualSpeechLM introduces a novel dual-token design by modeling USTokens as input and acoustic tokens as output via an integrated AcousticGPT, effectively accommodating different levels of information required for understanding and generation tasks.
@@
Next, the predicted USToken sequence $\mathbf{U^{tar}}$ is passed to AcousticGPT, which autoregressively generates acoustic tokens. They are then converted into speech via a Speech Decoder. For both understanding and generation, we adopt text LLM as backbone and fine-tune it using parameter-efficient Low-Rank Adaptation (LoRA) (Hu et al. 2022). Within the unified framework, we integrate AcousticGPT into text LLM to enable joint end-to-end training. This allows DualSpeechLM to simultaneously capture high-level semantics for understanding and fine-grained acoustic details for generation, which we refer to as dual-token modeling. Notably, the acoustic tokens and Speech Decoder are modular and can be implemented using any open-source codec model.
@@
Inspired by AudioLM (Borsos et al. 2023), we design an AcousticGPT module to autoregressively generate acoustic tokens conditioned on semantic hidden states and speaker embeddings. Built on a GPT-style transformer, the model consists of six causal blocks, each containing a causal self-attention layer followed by a Multi-Layer Perceptron (MLP). The multi-head self-attention captures dependencies under causal constraints. In addition, each block includes residual connections and layer normalization to enhance training stability and model performance. Further details are provided in Appendix B.
@@
$\mathcal{L}_{\text{acoustic}}=-\sum_{t=1}^{L}\log p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{E};\theta),$ (5)
where $\mathbf{S}\in\{\mathbf{S}_{p},\mathbf{U^{tar}},[\mathbf{S}_{p};\mathbf{U^{tar}}]\}$ is selected via CoC, $\mathbf{E}$ is the speaker embedding, and $\mathbf{A^{tar}}_{t}$ is the predicted acoustic token at time $t$. $p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{E};\theta)$ is the predicted probability of generating $\mathbf{A^{tar}}_{t}$, given $\{\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{E}\}$. Finally, the overall generation loss combines semantic and acoustic objectives as:
@@
We train the USTokenizer using multiple speech understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). They are based on LibriSpeech (Panayotov et al. 2015), IEMOCAP (Busso et al. 2008) and SQA dataset from SALMONN (Tang et al. 2023), respectively. In the SQA dataset, both questions and answers were automatically generated based on LibriSpeech transcripts using ChatGPT, forming spoken-question–text-answer pairs.
DualSpeechLM is trained on eight speech understanding and generation tasks using approximately 4,500 hours of speech data. For understanding, we use LibriSpeech for ASR task, the English-to-German (En2De) subset of CoVoST2 (Wang, Wu, and Pino 2020) for Speech-to-Text Translation (S2TT) task, IEMOCAP (Busso et al. 2008) for the SER task, and the SQA dataset (Tang et al. 2023) constructed from LibriSpeech using ChatGPT. For generation, we train Text-to-Speech (TTS) on LibriTTS-R (Koizumi et al. 2023), Text-to-Speech Translation (T2ST) on Spanish-to-English (Es–En) and French-to-English (Fr–En) subsets of CVSS (Jia et al. 2022), and Voice Conversion (VC) on a 1,000-hour subset of LibriHeavy (Kang et al. 2023) with evaluation on 400 pairs from VCTK (Yamagishi et al. 2019). Additionally, the Speech Conversation (SC) task is built by synthesizing speech from the SQA dataset using the Volcengine TTS API. Please see Appendix C for dataset details.
@@
In experiments, we compare four variants under the same configurations: (1) Baseline-Acoustic, where both input and output are acoustic token from WavTokenizer (Ji et al. 2024); (2) Baseline-Semantic, where both input and output are HuBERT token (Hsu et al. 2021); (3) DualSpeechLM-Hubert, using HuBERT token as input, acoustic token from WavTokenizer as output; and (4) DualSpeechLM-USToken, using our USToken as input, acoustic token from WavTokenizer as output. Notably, the Baseline-Acoustic model requires 160k steps to converge under multi-task settings, while others converge in 60k steps. Further details are provided in the Appendix D.

### v3

@@ Abstract
Extending pre-trained text Large Language Models (LLMs)’s speech understanding or generation abilities by introducing various effective speech tokens has attracted great attention in the speech research community. However, building a unified speech understanding and generation model still faces the following challenges: (1) Due to the huge modality gap between speech and text tokens, extending text LLMs to unified speech LLMs relies on large-scale paired data for fine-tuning, and (2) Generation and understanding tasks prefer information at different levels, e.g., generation benefits from detailed acoustic features, while understanding favors high-level semantics. This divergence leads to difficult performance optimization in one unified model. To solve these challenges, in this paper, we present two key insights in speech tokenization and speech language modeling. Specifically, we first propose an Understanding-driven Speech Tokenizer (USTokenizer), which extracts high-level semantic information essential for accomplishing understanding tasks using text LLMs. In this way, USToken enjoys better modality commonality with text, which reduces the difficulty of modality alignment in adapting text LLMs to speech LLMs. Secondly, we present DualSpeechLM, a dual-token modeling framework that concurrently models USToken as input and acoustic token as output within a unified, end-to-end framework, seamlessly integrating speech understanding and generation capabilities. Furthermore, we propose a novel semantic supervision loss and a Chain-of-Condition (CoC) strategy to stabilize model training and enhance speech generation performance. Experimental results demonstrate that our proposed approach effectively fosters a complementary relationship between understanding and generation tasks, highlighting the promising strategy of mutually enhancing both tasks in one unified model. [^1] Code and demo: https://github.com/lavendery/UUG.
@@ Introduction
Recent advancements in autoregressive large language models (LLMs) have demonstrated excellent performance in the natural language processing community (achiam2023gpt; touvron2023llama; team2023gemini). Leveraging the powerful foundations of text LLMs, recent advancements have led to emergence of speech LLMs that possess speech understanding and generation capabilities. In addition to fine-tuning textual LLMs to separately perform speech understanding tasks (gong2023joint; tang2023salmonn; chu2023qwenaudio; wang2024blsp; wang2025unisep) and generation tasks (wang2023neural; anastassiou2024seed; kim2024clam; wang2024speechx; yang2024uniaudio; yang2025simplespeech; jia2025ditar), developing unified models that excel at both capabilities has been explored in recent years (zhang2023speechgpt; xie2024miniomni2; fu2024vita; nguyen2025spirit; defossez2024moshi; xu2025qwen2). However, several limitations still remain.
@@
First, adapting pre-trained text LLMs to unified speech LLMs still relies heavily on large-scale paired speech-text data (zhang2023speechgpt; defossez2024moshi; xie2024miniomni2; xu2025qwen2). For example, SpeechGPT (zhang2023speechgpt) and SpiritLM (nguyen2025spirit) require approximately 70K and 570K hours of paired data, respectively. This dependence stems from the substantial modality gap between speech and text, which hinders capability transfer. Second, existing speech LLMs struggle to meet the distinct informational needs of understanding and generation. As shown in Figure 1 (a) Left and (b) Left, the Baseline model—trained solely with acoustic tokens—exhibits a contradiction between tasks, i.e., improving one often leads to degradation of the other, highlighting its inability to balance both tasks effectively. In fact, generation tasks demand rich acoustic details (e.g., prosody, emotion, speaker traits) for high-fidelity synthesis (zeghidour2021soundstream; defossez2022high; kumar2023high; wang2023neural), which acoustic tokens capture well but lack high-level semantics (shi2024espnet; dhawan2024codec; anastassiou2024seed; 10447929). Conversely, understanding tasks benefit from semantic features (borsos2023audiolm; rubenstein2023audiopalm; maiti2024voxtlm), but semantic tokens inevitably compromise the acoustic details needed for natural speech generation.
@@
To solve these problems, we propose two key insights from the perspectives of speech tokenization and language modeling. First, we present an Understanding-driven Speech Tokenizer (USTokenizer) that can extract high-level semantic features, which are critical for understanding tasks. Unlike prior methods that rely on only self-supervised learning (SSL) representation quantization (hsu2021hubert; chen2022wavlm) or automatic speech recognition (ASR)-based objectives (du2024cosyvoice; zeng2024glm4) to capture semantics, our approach directly aligns speech tokenizer with the semantic understanding capabilities of text LLMs. This leads to USTokens that have inherently better alignment with text modality, significantly easing modality alignment when adapting text LLMs to speech LLMs. Secondly, building on USTokenizer, we introduce DualSpeechLM, a novel dual speech token modeling framework that effectively handles the distinct informational needs of understanding and generation within a unified end-to-end architecture. Unlike conventional approaches that use the same token as input and output of LLM, our model separates them by using USTokens as input and acoustic tokens as output. Specifically, the USTokenizer provides high-level semantic information to boost understanding tasks, while an AcousticGPT module restores fine-grained acoustic details, enabling diverse and realistic speech generation within a unified end-to-end framework. In addition, we introduce a semantic supervision loss and a Chain of Condition (CoC) strategy to stabilize training and further improve the performance of the unified framework. In summary, our contributions are as follows:
@@
The success of autoregressive language models (achiam2023gpt; touvron2023llama; team2023gemini) has spurred progress in speech LLMs (chu2023qwenaudio; wang2024speechx; wang2024freeze; defossez2024moshi), where speech tokenizers are essential for converting continuous signals into discrete tokens (yangalmtokenizer). Speech tokenizers are typically categorized as acoustic or semantic (borsos2023audiolm; parkerscaling; yanguniaudio15). Acoustic tokens, optimized for signal reconstruction (zeghidour2021soundstream; defossez2022high; yang2023hifi; wang2023neural), capture detailed acoustic features beneficial for generation, but perform poorly on understanding tasks like ASR (shi2024espnet; dhawan2024codec; anastassiou2024seed; 10447929). Previous semantic tokenizers were trained in two ways: (1) applying clustering (borsos2023audiolm; zhang2023speechgpt; shi2023multi) or vector quantization (VQ) (huang2024repcodec) to representations of SSL models  (hsu2021hubert; chen2022wavlm), and (2) applying a VQ layer to the intermediate layer of ASR models  (du2024cosyvoice; zeng2024glm4; du2024cosyvoice2). Although these semantic tokenizers have shown benefits for understanding tasks (borsos2023audiolm; rubenstein2023audiopalm; maiti2024voxtlm), they do not explicitly consider the alignment to the modality of text LLM (li2025baichuanaudio). Furthermore, multi-codebook designs (zhang2023speechtokenizer; defossez2024moshi; yangalmtokenizer) aim to capture semantics and acoustics in different codebooks jointly, but often introduce complexity when integrated with LLMs. In this work, we present a single VQ-codebook USTokenizer, which not only incorporates high-level semantic information but also achieves better alignment between speech and text modality when applied to LLMs.
@@
Recent advances in speech LLMs have explored unified understanding and generation (zhang2023speechgpt; pan2024auto; defossez2024moshi; nguyen2025spirit). Some speech LLMs (yang2024uniaudio; shi2025balancing) adopt acoustic tokens to ensure high-fidelity speech synthesis. However, such tokens usually lead to degraded performance in understanding tasks, particularly in low-resource scenarios. By contrast, semantic tokens perform better in understanding tasks. Yet, their lack of acoustic details often results in reduced generation quality. To compensate for generation, existing works (polyak2021speech; zhang2023speechgpt; du2024cosyvoice; nguyen2025spirit) introduce additional components such as diffusion models (ho2020denoising) or flow matching (lipman2022flow), to convert speech tokens into a Mel spectrogram, and then a HiFi-GAN (kong2020hifi-gan) vocoder is used to synthesize waveform with the Mel spectrogram as input. These multi-stage pipelines increase complexity and risk error accumulation (jia2025ditar). To address these limitations, we propose DualSpeechLM, an end-to-end dual-token modeling framework that explicitly models USTokens as input for understanding and acoustic tokens as output for generation, effectively achieving distinct informational requirements.
@@
The Encoder converts Whisper features $\mathbf{E}$ into representations $\mathbf{H}$ using a 2× downsampling convolution followed by two residual convolutional blocks, then discretized by VQ. Symmetrically, the Decoder reconstructs $\mathbf{\hat{E}}$ from quantized vectors $\mathbf{\hat{H}}$. Both Encoder and Decoder adopt similar residual convolutional blocks with RepCodec (huang2024repcodec). We then compute the reconstruction loss via mean squared error (MSE) between $\mathbf{E}$ and $\mathbf{\hat{E}}$, formulated as:
@@
Figure 3: DualSpeechLM’s dual-token modeling paradigm. The left illustrates the baseline pipeline treating LLM input/output as identical tokens. In contrast, our DualSpeechLM (right) incorporates an Acoustic GPT module into the text LLM module for joint training, separately processing USToken inputs and acoustic token outputs through distinct modeling paths, effectively capturing the different levels of information required for both generation and understanding tasks.
@@
SpeechGPT  (zhang2023speechgpt) LLaMA-7B D(Hubert) Full FT 70K 42.73 78.54 1.07 - 3.58/40
SpiritLM  (nguyen2025spirit) LLaMA-7B D(Hubert) Full FT 570K 6.0 11.0 - - -
Mini-Omni2  (xie2024miniomni2) Qwen2-0.5B C(Whisper) Full FT 9K 4.8 9.8 - - -
VITA  (fu2024vita) Mixtral-8x7B C(Mel) Full FT 20K 8.14 18.41 - - 7.62/70
Moshi  (defossez2024moshi) Helium-7B D(Mimicodec) Full FT 7M 5.7 - - - -
Qwen2.5-Omni  (xu2025qwen2) Qwen2.5-7B C(Whisper) Full FT $\star$ 1.8 3.4 30.2 60.03 4.28/60
@@
The VQ module discretizes continuous feature representations $\mathbf{H}$ by mapping them to a codebook of learned vectors. In this work, we use a single VQ layer to quantize the feature $\mathbf{H}$ into discrete tokens $\mathbf{Z}$. Following previous works (huang2024repcodec), we add a commitment loss $\mathcal{L}_{\text{commit}}$ to ensure training stability, with more details provided in Appendix B.
@@
As shown in Figure 3, unlike prior Speech LLMs that use the same token for both input and output (yang2024uniaudio; wang2024speechx; shi2025balancing), DualSpeechLM introduces a novel dual-token design by modeling USTokens as input and acoustic tokens as output via an integrated AcousticGPT, effectively accommodating different levels of information required for understanding and generation tasks.
@@
Next, the predicted USToken sequence $\mathbf{U^{tar}}$ is passed to AcousticGPT, which autoregressively generates acoustic tokens. They are then converted into speech via a Speech Decoder. For both understanding and generation, we adopt text LLM as backbone and fine-tune it using parameter-efficient Low-Rank Adaptation (LoRA) (hu2022lora). Within the unified framework, we integrate AcousticGPT into text LLM to enable joint end-to-end training. This allows DualSpeechLM to simultaneously capture high-level semantics for understanding and fine-grained acoustic details for generation, which we refer to as dual-token modeling. Notably, the acoustic tokens and Speech Decoder are modular and can be implemented using any open-source codec model.
@@
Inspired by AudioLM (borsos2023audiolm), we design an AcousticGPT module to autoregressively generate acoustic tokens conditioned on semantic hidden states and speaker embeddings. Built on a GPT-style transformer, the model consists of six causal blocks, each containing a causal self-attention layer followed by a Multi-Layer Perceptron (MLP). The multi-head self-attention captures dependencies under causal constraints. In addition, each block includes residual connections and layer normalization to enhance training stability and model performance. Further details are provided in Appendix B.
@@
$\mathcal{L}_{\text{acoustic}}=-\sum_{t=1}^{L}\log p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk};\theta),$ (5)
where $\mathbf{S}\in\{\mathbf{S}_{p},\mathbf{U^{tar}},[\mathbf{S}_{p};\mathbf{U^{tar}}]\}$ is selected via CoC, $\mathbf{Spk}$ is the speaker embedding, and $\mathbf{A^{tar}}_{t}$ is the predicted acoustic token at time $t$. $p(\mathbf{A^{tar}}_{t}|\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk};\theta)$ is the predicted probability of generating $\mathbf{A^{tar}}_{t}$, given $\{\mathbf{A^{tar}}_{<t},\mathbf{S},\mathbf{Spk}\}$. Finally, the overall generation loss combines semantic and acoustic objectives as:
@@
We train the USTokenizer using multiple speech understanding tasks, including Automatic Speech Recognition (ASR), Speech Emotion Recognition (SER), and Speech Question Answering (SQA). They are based on LibriSpeech (7178964), IEMOCAP (busso2008iemocap) and SQA dataset from SALMONN (tang2023salmonn), respectively. In the SQA dataset, both questions and answers were automatically generated based on LibriSpeech transcripts using ChatGPT, forming spoken-question–text-answer pairs.
DualSpeechLM is trained on eight speech understanding and generation tasks using approximately 4,500 hours of speech data. For understanding, we use LibriSpeech for ASR task, the English-to-German (En2De) subset of CoVoST2 (wang2020covost) for Speech-to-Text Translation (S2TT) task, IEMOCAP (busso2008iemocap) for the SER task, and the SQA dataset (tang2023salmonn) constructed from LibriSpeech using ChatGPT. For generation, we train Text-to-Speech (TTS) on LibriTTS-R (koizumi2023librittsr), Text-to-Speech Translation (T2ST) on Spanish-to-English (Es–En) and French-to-English (Fr–En) subsets of CVSS (jia2022cvss), and Voice Conversion (VC) on a 1,000-hour subset of LibriHeavy (kang2023libriheavy) with evaluation on 400 pairs from VCTK (Yamagishi2019CSTRVC). Additionally, the Speech Conversation (SC) task is built by synthesizing speech from the SQA dataset using the Volcengine TTS API. Please see Appendix C for dataset details.
@@
In experiments, we compare four variants under the same configurations: (1) Baseline-Acoustic, where both input and output are acoustic token from WavTokenizer (ji2024wavtokenizer); (2) Baseline-Semantic, where both input and output are HuBERT token (hsu2021hubert); (3) DualSpeechLM-Hubert, using HuBERT token as input, acoustic token from WavTokenizer as output; and (4) DualSpeechLM-USToken, using our USToken as input, acoustic token from WavTokenizer as output. Notably, the Baseline-Acoustic model requires 160k steps to converge under multi-task settings, while others converge in 60k steps. Further details are provided in the Appendix D.